A 4.85 g sample of anhydrous sodium sulfate is dissolved in water and the solution made up to 250 cm3 in a volumetric flask.
What is the concentration in mol dm−3 of sodium sulfate in the solution?
1
Expert's answer
2020-07-07T14:19:21-0400

Concentration= amount of solute/volume im dm3

Solution in dm3 = 250/1000

=0.25dm3

now concentration in g/dm3=4.85/0.25

=19.4g/dm3

Molecular mass of sodium sulfate is 142.04

Now concentration in mol/dm3 becomes,

19.4/142.04

=0.13mol/dm3